OpenStudy (gator95):

What number should be added to 17 to make the sum 0?

OpenStudy (hayisforhorses):

17+x=0 -17 -17 x=-17

OpenStudy (avirllavinge2017):

17 is what you call a "positive" numeral. To make a sum of 0 you would have to switch that "positive" numeral around to what is called a "negative" numeral. Negative numerals are always minus that number Therefore , @hayisforhorses gave a good answer but didn't explain too much to what he or she got , so he or she didn't give enough information. However , Lets start doing shall we! so we have our number 17 , we need to make a sum of 0. so you are prob saying how? right? you cant "add" anything to positive 17 , so you can add to the 0 but not with a positive 17 , you would have to subtract that 0 with the negative 17 , but all you really do to sum it up is to just switch the sides!! since you cant add the positive 17 , you have to make that 17 a negative to make a sum up to 0, so 17+x-0? now switch!! X/ -17 < answer

OpenStudy (avirllavinge2017):

@ILovePuppiesLol is this a good way of explaining? too much?

ILovePuppiesLol (ilovepuppieslol):

you are all good :) just keep doing what ur doing! @AvirlLavinge2017